# Composite likelihood inference for the Poisson log-normal model

### v1

@@ 4.1 Determining the blocks for composite likelihood inference
To get an estimate of each entry of the covariance matrix $\Sigma$, it is sufficient that each pair of distinct species $(j,j^{\prime})$ ($1\leq j<j^{\prime}\leq p$) appears at least once in a same block $\mathcal{C}_{b}$ ($1\leq b\leq C$). Hence, it is unnecessary to explore every possible combinations of blocks with size $k$, so $C\leq{{p}\choose{k}}$. On the other hand, because there are $p(p-1)/2$ pairs of species and because each block contains $k(k-1)/2$ pairs (giving a total of $Ck(k-1)/2$ pairs), we need that $C\geq p(p-1)/[k(k-1)]$. Remark that $k=2$ is a trivial case as each blocks contributes to estimate one single covariance parameter so $C=p(p-1)/2$.
@@ 4.2 Simulation study
[…]
@@
CL2
CL3
CL5
CL7
@@
Computational time (s)
CL2222 CL3333 CL5555 CL7777
$p=10$ 44 42 216 149
$p=30$ 4061 8936 592 3196
$p=50$ 10769 8744 4249
Number of iterations
CL2222 CL3333 CL5555 CL7777
$p=10$ 124 115 107 101
$p=30$ 299 256 243 216
$p=50$ 522 443 123
Number of blocks
CL2222 CL3333 CL5555 CL7777
$p=10$ 45 17 6 3
$p=30$ 435 159 60 34
$p=50$ 448 166 93
@@ 4.3 Fish abundances in the Barents sea
Reduced dataset: $p=7$ species
@@
FL
CL2
CL3
CL5
Figure 5: (Composite) log-likelihood (left) and effective sample size (right) along iterations for the different algorithms. From top to bottom: FL, CL2, CL3, CL5. Left: one dot for each iteration, right: one boxplot for each iteration (one value for each block and each site).
@@
Figure 6: Comparison of the estimates with the different algorithms. Left: regression coefficients $\widehat{\beta}_{hj}$, center: covariance parameters $\widehat{\Sigma}_{jk}$, right: estimated variances of the regression coefficients $\widehat{\beta}_{hj}$. $x$-axis = full likelihood inference (FL), $y$-axis: likelihood inference (FL = red triangle $\triangle$ = reference), composite likelihood inference (CL2 = green plus $+$, CL3 = blue times symbol $\times$, CL5 = cyan diamond $\diamond$) and variational EM (VEM = black circles $\circ$’).
@@
Full dataset: $p=30$ species
@@
The center and right panels of Figure 8 presents the estimate of the latent covariance matrix $\Sigma$ (and of the associated correlation matrix $\text{cor}({\boldsymbol{\Sigma}})$), where only significant terms are displayed. This results in a fairly sparse pattern, indicating that only a small fraction of species pairs (41 among 435) have correlated abundance variations, once accounted for the environmental effects.
@@
Figure 8: Colormap of the parameter estimates. Left: regression coefficients $\widehat{\beta}_{\ell j}$, center: covariance parameters $\widehat{\Sigma}_{jk}$, right: corresponding correlations $\widehat{\rho}_{jk}=\widehat{\Sigma}_{jk}/\sqrt{\widehat{\Sigma}_{jj}\widehat{\Sigma}_{kk}}$. Blank cells correspond to non-significant estimates, red cells (marked with ’$+$’) to positive estimates and blue cells (marked with ’--’) to negative estimates.
@@ Acknowledgements
The authors are grateful to the INRAE MIGALE bioinformatics facility (MIGALE, INRAE, 2020. Migale bioinformatics Facility, doi: 10.15454/1.5572390655343293E12) for providing computing resources. The first author has been partly funded by the European Union (ERC-2022-SYG-OCEAN-101071601). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ A.1 Proofs
Proof of Proposition 1
As $\theta^{(h+1)}$ is the maximizer of $\sum_{b=1}^{B}\lambda_{b}\mathbb{E}_{\theta^{(h)}}\left[\log p_{\theta}({\mathbf{Y}}^{(b)},{\mathbf{Z}}^{(b)})\;|\;{\mathbf{Y}}^{(b)}\right]$, we have that
00 0 $\leq\sum_{b=1}^{B}\lambda_{b}\mathbb{E}_{\theta^{(h)}}\left[\log p_{\theta^{(h+1)}}({\mathbf{Y}}^{(b)},{\mathbf{Z}}^{(b)})\;|\;{\mathbf{Y}}^{(b)}\right]-\sum_{b=1}^{B}\lambda_{b}\mathbb{E}_{\theta^{(h)}}\left[\log p_{\theta^{(h)}}({\mathbf{Y}}^{(b)},{\mathbf{Z}}^{(b)})\;|\;{\mathbf{Y}}^{(b)}\right]$
$=\sum_{b=1}^{B}\lambda_{b}\mathbb{E}_{\theta^{(h)}}\left[\log\frac{p_{\theta^{(h+1)}}({\mathbf{Y}}^{(b)},{\mathbf{Z}}^{(b)})}{p_{\theta^{(h)}}({\mathbf{Y}}^{(b)},{\mathbf{Z}}^{(b)})}\;|\;{\mathbf{Y}}^{(b)}\right]\leq\sum_{b=1}^{B}\lambda_{b}\log\mathbb{E}_{\theta^{(h)}}\left[\frac{p_{\theta^{(h+1)}}({\mathbf{Y}}^{(b)},{\mathbf{Z}}^{(b)})}{p_{\theta^{(h)}}({\mathbf{Y}}^{(b)},{\mathbf{Z}}^{(b)})}\;|\;{\mathbf{Y}}^{(b)}\right]$
@@
$\frac{p_{{\boldsymbol{\theta}}^{(h)}}({\mathbf{Y}}_{i},{\mathbf{v}})^{2}}{\varphi({\mathbf{v}};m,S)}$ $=\frac{|B|^{\nicefrac{{1}}{{2}}}}{(2\pi)^{\nicefrac{{p}}{{2}}}\left|\Sigma\right|}\exp\left\{-\frac{1}{2}{\mathbf{v}}^{\top}\left(2\Sigma^{-1}-S^{-1}\right){\mathbf{v}}\right\}$
@@ A.2 Update formulas (M step)
Update for ${\boldsymbol{B}}$.
For all $(i,j)\in\llbracket 1,n\rrbracket\times\llbracket 1,p\rrbracket$
@@
Update for ${\boldsymbol{\Sigma}}$.
For all $(i,j,k)\in\llbracket 1,n\rrbracket\times\llbracket 1,p\rrbracket\times\llbracket 1,p\rrbracket$,
@@ A.3 Update formulas (CL-M step)
Update for ${\boldsymbol{B}}$.
For all $(i,j,b)\in\llbracket 1,n\rrbracket\times\llbracket 1,p\rrbracket\times\llbracket 1,C\rrbracket$,
@@
If we denote $\mathcal{C}(j)=\{b\in\llbracket 1,C\rrbracket\;|\;j\in\mathcal{C}_{b}\}$ the set of indices of blocks that contain $j$, the gradient in ${\boldsymbol{\beta}}_{j}$ writes as
@@
Update for ${\boldsymbol{\Sigma}}$.
Unlike the full data framework we do not have access to an estimate of $\Sigma$ and should resort as well to a gradient based method. Denote $\Omega^{(b)}$ the inverse matrix of $\Sigma^{(b)}$ and for $(j,k)\in\llbracket 1,p\rrbracket^{2}$, $\mathcal{C}(j,k)=\{b\in\llbracket 1,C\rrbracket\;|\;j,k\in\mathcal{C}_{b}\}$
@@ A.4 Additional simulation results
Full likelihood inference
We present here additional results regarding the distribution of the full likelihood estimates of the regression coefficients $\beta_{ellj}$.
Figure 9: Distribution of the $p$-values of the Kolmogorov-Smirnov test for the distribution of the standardized estimates ${\widetilde{\beta}}_{\ell j}$ (over the $M=100$ simulations) as a function of the number of species ($p=5,\dots 10$) for full-likelihood inference (FL). Dotted red lines: $\alpha=5\%$ significance threshold after Bonferronni correction (i.e., $\alpha/(dp)$).
@@
$p=5$
$p=6$
$p=7$
$p=8$
$p=9$
$p=10$
Figure 10: qq-plots of the normalized regression coefficient ${\widetilde{\beta}}_{\ell j}$ defined in Equation (20) for the second simulated species for each of the $d=3$ covariates with full likelihood (FL) with $p=5\dots 10$ species. Same legend as Figure 3.
Composite likelihood inference
This section provides additional results regarding the distribution of the composite likelihood estimates of the regression coefficients $\beta_{ellj}$.
@@
FL
CL2
CL3
CL5
CL7

### v2

@@ 4.2 Determining the blocks for composite likelihood inference
To get an estimate of each entry of the covariance matrix $\Sigma$, it is sufficient that each pair of distinct species $(j,j^{\prime})$, $1\leq j<j^{\prime}\leq p$, appears at least once in a same block $\mathcal{C}_{b}$, $1\leq b\leq C$. Hence, it is unnecessary to explore every possible combinations of blocks with size $k$, so $C\leq{{p}\choose{k}}$. On the other hand, because there are $p(p-1)/2$ pairs of species and because each block contains $k(k-1)/2$ pairs (giving a total of $Ck(k-1)/2$ pairs), we need that $C\geq p(p-1)/[k(k-1)]$. Remark that $k=2$ is a trivial case as each blocks contributes to estimate one single covariance parameter so $C=p(p-1)/2$.
@@ 4.3.1 Simulation design
For all ISEM algorithms, we used a linearly increasing number of particles along the iterations: at iteration $h$, we used $N^{(h)}=hN^{(0)}$ particles. We considered $N^{(0)}=50,100$ and $200$ initial particles. We did not observe any significant difference, except in the situations (not shown) that are actually doomed to fail with all considered numbers of particles due to the aforementioned limitation of IS when the dimension $p$ or $k$ grows. We report hereafter the results obtained with $N^{(0)}=200$ initial particles. We let the algorithms run at most 1000 iterations and used a lag of 50 steps for the stopping criterion. Regarding the mixture proposal distribution defined in Equation (20), we used the mixture proportion $\alpha=0.9$.
@@ 4.3.2 Simulation results
CL2 Refer to caption Refer to caption Refer to caption
CL3 Refer to caption Refer to caption Refer to caption
CL5 Refer to caption Refer to caption Refer to caption
CL7 Refer to caption Refer to caption Refer to caption
VEM Refer to caption Refer to caption Refer to caption
@@
Computational time (s) Number of iterations Number of blocks
Algorithm CL2222 CL3333 CL5555 CL7777 VEM CL2222 CL3333 CL5555 CL7777 CL2222 CL3333 CL5555 CL7777
$p=10$ 44 42 216 149 10 124 115 107 101 45 17 6 3
$p=30$ 4061 8936 592 3196 18 299 256 243 216 435 159 60 34
$p=50$ 10769 8744 4249 34 522 443 123 448 166 93
Comparison with VEM estimates.
[…]
Unfortunately, Figure 2 shows that the discrepancy between the distribution of the standardized statistics $\widetilde{\beta}_{k\ell}$ and a standard normal is poor (normality is rejected in a large proportion of simulations). Figure 3 (last row) shows that this poor fit is mostly due to an over-estimation of the variance, resulting in too wide confidence intervals ans poor statistical power for hypothesis testing.
@@ 4.4 Fish abundances in the Barents sea
4.4.1 Reduced datas et: $p=7$ species
@@
FL Refer to caption Refer to caption
CL2 Refer to caption Refer to caption
CL3 Refer to caption Refer to caption
CL5 Refer to caption Refer to caption
Figure 5: (Composite) log-likelihood (left) and effective sample size (right) along iterations for the different algorithms and Barents sea reduced data set ($p=7$ species). From top to bottom: FL, CL2, CL3, CL5. Left: one dot for each iteration, right: one boxplot for each iteration (one value for each block and each site).
@@
[…]
@@
4.4.2 Full data set: $p=30$ species
@@
To further inquire difference between the VEM and CL alogithms, we checked the proportion of VEM estimates covered by the CL confidence intervals. We observed that only 4 variational estimates of the regression coefficients $\beta_{\ell j}$ (out of $q\times p=150$) felt outside the CL5 composite likelihood confidence interval and that all variational estimates of the variance parameters $\sigma_{kj}$ where inside the CL5 confidence intervals. To this respect, only a very small fraction of the VEM estimates turned out to significantly different from the corresponding CL5 estimates. Still, Table 4 shows that, when such a discrepancy occurs, it may lead to substantially different estimations of the regression parameters.
Table 4: Comparison of the VEM and CL5 estimates when the VEM estimate does not meet the CL5 95% confidence interval for the Barents sea full data set ($p=30$ species).
Regression parameter $\beta_{1,1}$ $\beta_{2,2}$ $\beta_{2,2}$ $\beta_{1,11}$
VEM -1.23 2.46 2.09 -0.94
CL5 -6.33 5.39 3.10 -1.93
Model selection.
As reminded at the end of Section 2.3, a BIC criterion can be associated with composite likelihood inference. Combining the four available covariate, we fitted the $2^{4}$ possible models (from Model 1: intercept only, Model 2: intercept plus Latitude, Model 3: intercept plus Longitude, up to Model 16: intercept plus the four covariates) to the full Barents data set. For each model, we estimated the dimension $\text{dim}({\boldsymbol{\theta}})$ defined in (14) as $\widehat{\text{dim}}({\boldsymbol{\theta}})=\mathrm{tr}[\widehat{H}_{n}({\boldsymbol{\theta}})\widehat{G}_{n}({\boldsymbol{\theta}})^{-1}]$ where $\widehat{H}_{n}({\boldsymbol{\theta}})$ and $\widehat{J}_{n}({\boldsymbol{\theta}})$ are given in (12) and $\widehat{G}_{n}({\boldsymbol{\theta}})=\widehat{H}_{n}({\boldsymbol{\theta}})\widehat{J}_{n}({\boldsymbol{\theta}})^{-1}\widehat{H}_{n}({\boldsymbol{\theta}})$.
The left panel of Figure 8 gives the values of the composite log-likelihoods obtained with the CL5 and CL7 algorithms. We subtracted the composite log-likelihood for Model 1 to emphasize that $c\ell_{\widehat{{\boldsymbol{\theta}}}}(bY)$ does not increase at the same speed, depending on the block size $k=5$ or $7$. This is consistent with the notion of adaptive dimension $\dim({\boldsymbol{\theta}})$ appearing in the BIC derived by Gao and Song (2010) (see (14)).
The center panel of Figure 8 shows that, although the estimate $\widehat{\text{dim}}({\boldsymbol{\theta}})$ of this adaptive dimension is a combination of the two Monte-Carlo estimates $\widehat{H}_{n}({\boldsymbol{\theta}})$ and $\widehat{J}_{n}({\boldsymbol{\theta}})$, it remains almost constant among models involving the same number of variables.
$c\ell_{\widehat{{\boldsymbol{\theta}}}_{k}}$ $\text{dim}(\widehat{{\boldsymbol{\theta}}}_{k})$ $BIC(k)$
Refer to caption Refer to caption Refer to caption
Figure 8: Model selection among the 16 possible models ($x$-axis) for the Barents sea full data set ($p=30$ species) with the CL5 and CL7 algorithms. Dashed vertical lines: limits between models involving $d=1,\dots 5$ covariates (including the intercept). Left: composite likelihood (subtracting the composite likelihood for Model 1); Center: dimension of the parameter, as defined in Equation (14); Right: BIC criterion (subtracting the composite likelihood for Model 1). CL5 algorithm: red triangle [$\triangle$], CL7: blue ’times’ sign [$\times$] (the best model within each dimension is circled for each algorithm).
The right panel of Figure 8 gives the final BIC criteria for the 16 models and the two algorithms CL5 and CL7 (still subtracting the composite log-likelihood of Model 1). When considering models with same number of covariates, the BIC criterion does select the same model with one covariate (Latitude), two covariates (Latitude + Depth) and four covariates for the CL5 and CL7 algorithm. However the model Latitude + Depth + Temperature is selected with CL5, whereas CL7 yields the model Latitude + Depth + Longitude. This differences are likely to result from the strong correlation between the Temperature and both the Latitude and the Longitude that is observed in the data set.
Overall, for both algorithms, the BIC criterion opts for the full model (Model 16), suggesting that all covariates should be kept in the model. This is consistent with the fact that, when considering a separate univariate PLN model for each of the $p=30$ species, each of the four covariates turns out to be significant for at least one of the species, even after Bonferroni correction (not shown). We will therefore focus on the interpretation of the full model, in the next section.
@@
The center and right panels of Figure 9 presents the estimate of the latent covariance matrix ${\boldsymbol{\Sigma}}$ (and of the associated correlation matrix $\text{cor}({\boldsymbol{\Sigma}})$), where only significant terms are displayed. This results in a fairly sparse pattern, indicating that only a small fraction of species pairs (41 among 435) have correlated abundance variations, once accounted for the environmental effects.
@@
Figure 9: Colormap of the parameter estimates the Barents sea full data set ($p=30$ species). Left: regression coefficients $\widehat{\beta}_{\ell j}$, center: covariance parameters $\widehat{\Sigma}_{jk}$, right: corresponding correlations $\widehat{\rho}_{jk}=\widehat{\Sigma}_{jk}/\sqrt{\widehat{\Sigma}_{jj}\widehat{\Sigma}_{kk}}$. Blank cells correspond to non-significant estimates, red cells (marked with ’$+$’) to positive estimates and blue cells (marked with ’--’) to negative estimates.
@@ Acknowledgements
The authors thank Julien Chiquet (INRAE MIA-Paris-Saclay, France) and Mahendra Mariadassou (INRAE MaIAGE, France) for helpful discussions and advice. The authors are grateful to the INRAE MIGALE bioinformatics facility (MIGALE, INRAE, 2020. Migale bioinformatics Facility, doi: 10.15454/1.5572390655343293E12) for providing computing resources. The first author has been partly funded by the European Union (ERC-2022-SYG-OCEAN-101071601). Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ A.1 Proofs
A.1.1 Proof of Proposition 1
As $\theta^{(h+1)}$ is the maximizer of $\sum_{b=1}^{B}\lambda_{b}\mathbb{E}_{\theta^{(h)}}\mathopen{}\mathclose{{}\left[\log p_{\theta}({\mathbf{Y}}^{(b)},{\mathbf{Z}}^{(b)})\;|\;{\mathbf{Y}}^{(b)}}\right]$, we have that
00 0 $\leq\sum_{b=1}^{B}\lambda_{b}\mathbb{E}_{\theta^{(h)}}\mathopen{}\mathclose{{}\left[\log p_{\theta^{(h+1)}}({\mathbf{Y}}^{(b)},{\mathbf{Z}}^{(b)})\;|\;{\mathbf{Y}}^{(b)}}\right]-\sum_{b=1}^{B}\lambda_{b}\mathbb{E}_{\theta^{(h)}}\mathopen{}\mathclose{{}\left[\log p_{\theta^{(h)}}({\mathbf{Y}}^{(b)},{\mathbf{Z}}^{(b)})\;|\;{\mathbf{Y}}^{(b)}}\right]$
$=\sum_{b=1}^{B}\lambda_{b}\mathbb{E}_{\theta^{(h)}}\mathopen{}\mathclose{{}\left[\log\frac{p_{\theta^{(h+1)}}({\mathbf{Y}}^{(b)},{\mathbf{Z}}^{(b)})}{p_{\theta^{(h)}}({\mathbf{Y}}^{(b)},{\mathbf{Z}}^{(b)})}\;|\;{\mathbf{Y}}^{(b)}}\right]\leq\sum_{b=1}^{B}\lambda_{b}\log\mathbb{E}_{\theta^{(h)}}\mathopen{}\mathclose{{}\left[\frac{p_{\theta^{(h+1)}}({\mathbf{Y}}^{(b)},{\mathbf{Z}}^{(b)})}{p_{\theta^{(h)}}({\mathbf{Y}}^{(b)},{\mathbf{Z}}^{(b)})}\;|\;{\mathbf{Y}}^{(b)}}\right]$
@@ A.1.2 Proof of Proposition 3
$\frac{p_{{\boldsymbol{\theta}}^{(h)}}({\mathbf{Y}}_{i},{\mathbf{v}})^{2}}{\varphi({\mathbf{v}};m,S)}$ $=\frac{|S|^{\nicefrac{{1}}{{2}}}}{(2\pi)^{\nicefrac{{p}}{{2}}}\left|\Sigma\right|}\exp\left\{-\frac{1}{2}{\mathbf{v}}^{\top}\left(2\Sigma^{-1}-S^{-1}\right){\mathbf{v}}\right\}$
@@ A.2 Update formulas (M step)
Update for ${\boldsymbol{B}}$.   For all $(i,j)\in\{1,\ldots,n\}\times\{1,\ldots,p\}$
@@
Update for ${\boldsymbol{\Sigma}}$.   For all $(i,j,k)\in\{1,\ldots,n\}\times\{1,\ldots,p\}\times\{1,\ldots,p\}$,
@@ A.3 Update formulas (CL-M step)
Update for ${\boldsymbol{B}}$.   For all $(i,j,b)\in\{1,\ldots,n\}\times\{1,\ldots,p\}\times\{1,\ldots,C\}$,
@@
If we denote $\mathcal{C}(j)=\{b\in\{1,\ldots,C\}\;|\;j\in\mathcal{C}_{b}\}$ the set of indices of blocks that contain $j$, the gradient in ${\boldsymbol{\beta}}_{j}$ writes as
@@
Update for ${\boldsymbol{\Sigma}}$.   Unlike the full data framework we do not have access to an estimate of $\Sigma$ and should resort as well to a gradient based method. Denote $\Omega^{(b)}$ the inverse matrix of $\Sigma^{(b)}$ and for $(j,k)\in\{1,\ldots,p\}^{2}$, $\mathcal{C}(j,k)=\left\{b\in\{1,\ldots,C\}\;|\;j,k\in\mathcal{C}_{b}\right\}$
@@ A.4 Additional simulation results
A.4.1 Full likelihood inference
We present here additional results regarding the distribution of the full likelihood estimates of the regression coefficients $\beta_{\ell j}$.
Figure 10: Distribution of the $p$-values of the Kolmogorov-Smirnov test for the distribution of the standardized estimates ${\widetilde{\beta}}_{\ell j}$ over the $M=100$ simulations as a function of the number of species ($p=5,\dots,10$) for full-likelihood inference (FL). Dotted red lines: $\alpha=5\%$ significance threshold after Bonferronni correction (i.e., $\alpha/(dp)$).
@@ A.4.1 Full likelihood inference
$p=5$ Refer to caption Refer to caption Refer to caption
$p=6$ Refer to caption Refer to caption Refer to caption
$p=7$ Refer to caption Refer to caption Refer to caption
$p=8$ Refer to caption Refer to caption Refer to caption
$p=9$ Refer to caption Refer to caption Refer to caption
$p=10$ Refer to caption Refer to caption Refer to caption
Figure 11: qq-plots of the normalized regression coefficient ${\widetilde{\beta}}_{\ell j}$ defined in Equation (22) for the second simulated species for each of the $d=3$ covariates with full likelihood (FL) with $p=5,\dots,10$ species. Same legend as Figure 3.
A.4.2 Composite likelihood inference
This section provides additional results regarding the distribution of the composite likelihood estimates of the regression coefficients $\beta_{\ell j}$.
@@ A.4.2 Composite likelihood inference
FL Refer to caption Refer to caption Refer to caption
CL2 Refer to caption Refer to caption Refer to caption
CL3 Refer to caption Refer to caption Refer to caption
CL5 Refer to caption Refer to caption Refer to caption
CL7 Refer to caption Refer to caption Refer to caption